• Wrong document? Swap it for free
  • Written by students who passed
  • Immediately available after payment
  • Read online or as PDF
Sell
Where do you study
Your language
Document preview thumbnail
Preview 2 out of 7 pages
Exam (elaborations)

INDU 6331 Assignment 2 Advanced Quality Control

Document preview thumbnail
Preview 2 out of 7 pages

A mechatronic assembly is subjected to a final functional test. Suppose that defects occur at random in these assemblies, and that defects occur according to a Poisson distribution with parameter λ= 0.02 a. What is the probability that an assembly will have exactly one defect? λ=0.02 p ( x )= e −λ λ x x ! p( x=1)= e −0.02 0.021 1! =0.0196=1.96 There is a probability of 1.96% that an assembly will have exactly one defect b. What is the probability that an assembly will have one or more defects? p(x=1) = 1 – p(x=0) p( x≥1)=1− e −0.020.020 0! =1−0.98019=0.0198 The probability that an assembly will have one or more defects is 1.98% c. Suppose that you improve the process so that the occurrence rate of defects is cut in half to λ= 0.01. What effect that this has on the probability that an assembly will have one or more defects? P(x=1) = 1 – p(x=0) p( x≥1)=1− e −0.010.010 0! =1−0.99=0.00995 The probability that an assembly will have one or more defects considering λ= 0.01, is 0.995%. Cutting in half the rate of defects reduces the probability calculated to its half. 3.26 The probability distribution of x is f(x)= ke-x , 0≤ x ≤∞. Find the appropriate value of k. Find the mean and variance of x. This study source was downloaded by from CourseH on :11:11 GMT -05:00 INDU 6331 Assignment 2 Advanced Quality Control For this function the appropriate value of k is k= 1. With this value f(x) will describe an exponential distribution: f ( x )=k e −kx , k=1 The mean of this distribution is μ= 1 k = 1 1 =1 And the variance is σ 2= 1 k 2= 1 1 =1 3.33 A random sample of 50 units is drawn from a production process every half hour. The fraction of nonconforming product manufactured is 0.02. What is the probability that Ṗ≤ 0.04 if the fraction nonconforming really is 0.02? Binomial distribution n= 50 p= 0.02 Ṗ≤0.04 Ṗ= x n ≤0.04→ x 50 ≤ 0.04→x ≤2 Probability of finding 2 or less nonconformings in the sample


Document information

Uploaded on
June 25, 2023
Number of pages
7
Written in
2022/2023
Type
Exam (elaborations)
Contains
Questions & answers
$7.69

Wrong document? Swap it for free Within 14 days of purchase and before downloading, you can choose a different document. You can simply spend the amount again.
Written by students who passed
Immediately available after payment
Read online or as PDF

Seller avatar
Reputation scores are based on the amount of documents a seller has sold for a fee and the reviews they have received for those documents. There are three levels: Bronze, Silver and Gold. The better the reputation, the more your can rely on the quality of the sellers work.
ExamsRevision
4.4
(45)
Sold
216
Followers
49
Items
1541
Last sold
3 weeks ago




Why students choose Stuvia

Created by fellow students, verified by reviews

Quality you can trust: written by students who passed their tests and reviewed by others who've used these notes.

Didn't get what you expected? Choose another document

No worries! You can instantly pick a different document that better fits what you're looking for.

Pay as you like, start learning right away

No subscription, no commitments. Pay the way you're used to via credit card and download your PDF document instantly.

Student with book image

“Bought, downloaded, and aced it. It really can be that simple.”

Alisha Student

Working on your references?

Create accurate citations in APA, MLA and Harvard with our free citation generator.

Working on your references?

Frequently asked questions

Whoops! We can’t load your doc right now. Try again or contact support.